# Human-AI Collaborative Multi-modal Multi-rater Learning for Endometriosis Diagnosis

### v2

@@ 4.5 Human-AI Collaborative Multi-modal Multi-rater Ablation Study
In this subsection, we show an ablation study of different combinations of annotators to be used in the human-AI collaboration and model training. We also analyze the use of multi-modal data in the results.
@@
Table 2: Accuracy and AUROC performance analyses of HAICOMM and its variants. “R1”, “R2” and “R3” denote models trained with input annotations from annotators #1, #2 and #3, respectively. “HAIC” represents model trained with multi-rater labels inputs for human-AI collaborations. T1 and T2 Only w/ HAIC represent single-modality HAICOMM approaches. The best results for each column are in bold.
Models Accuracy AUROC
Labels from Rater #1 (R1) 0.6667 -
Labels from Rater #2 (R2) 0.7333 -
Labels from Rater #3 (R3) $0$
HAICOMM w/o HAIC 0.6333 0.5933
HAICOMM w/ R1 0.7667 0.7044
HAICOMM w/ R2 0.8000 0.8444
HAICOMM w/ R3 0.5667 0.5956
HAICOMM w/ R1,2 0.7667 0.8667
HAICOMM w/ R2,3 0.7333 0.8711
HAICOMM w/ R1,3 0.6333 0.6689
T1 Only w/ HAIC 0.6667 0.8067
T2 Only w/ HAIC 0.7667 0.8844
HAICOMM 0.8000 0.8865
4.6 Qualitative Analyses
We also conduct qualitative analyses about HAICOMM. In Figure 2, (a) and (b) are the input T1 and T2 MRIs, respectively. The table below shows the predictions by the three raters (Rater #1,#2,#3), then the predictions by SSR and ProMix trained with Rater #1’s labels and CROWDLAB’s labels (SSR w GT1, ProMix w GT1, SSR w CL GT, ProMix w CL GT). Next, we show SSR and ProMix trained with CROWDLAB’s labels and relying on human-AI collaborative classification (SSR w HAIC, ProMix w HAIC), followed by the result from our HAICOMM, and the ground truth label from surgical data. The case shows the proposed HAICOMM model can generate correct labels while other models cannot. For Figure 3, the case also shows that the proposed HAICOMM model can generate a correct label while most other methods cannot (only ProMix w GT1 and HAICOMM predict the surgical ground truth label correctly).
Figure 2: Qualitative example analysis of HAICOMM.
Figure 3: Another qualitative example analysis of HAICOMM.
@@ 5 Conclusion and Future Work
One potential limitation of our method is the dataset size. Currently, we are dedicated to collect more data from different clinical sources to expand the dataset. The use of such multiple clinical sources will require the exploration of domain adaptation techniques to enable a better flexibility of the method to work in multiple domains. Beyond this issue, the need for specific labellers for training and testing is another potential limitation. We plan to address this issue with the development of techniques that work with a variable set of labellers during training and testing. Another interesting direction is the collection of new datasets for other multi-modal multi-rater clinical problems to enable the evaluation of our HAICOMM in a different task.

### v3

@@ 5.2 Human-AI Collaborative Multi-modal Multi-rater Ablation Study
Table 5: Accuracy and AUROC performance analyses of HAICOMM and its variants on the test set. “R1”, “R2” and “R3” denote models trained with input annotations from annotators #1, #2 and #3, respectively. “HAIC” represents model trained with multi-rater labels inputs for human-AI collaborations. T1 and T2 Only w/ HAIC represent single-modality HAICOMM approaches. The best results for each column are in bold.
Models Accuracy AUROC
Labels from Rater #1 (R1) 0.67 -
Labels from Rater #2 (R2) 0.73 -
Labels from Rater #3 (R3) 0.70 -
HAICOMM w/o HAIC 0.63 0.59
HAICOMM w/ R1 0.77 0.70
HAICOMM w/ R2 0.80 0.84
HAICOMM w/ R3 0.57 0.60
HAICOMM w/ R1,2 0.77 0.87
HAICOMM w/ R2,3 0.73 0.87
HAICOMM w/ R1,3 0.63 0.67
T1 Only w/ HAIC 0.67 0.81
T2 Only w/ HAIC 0.77 0.88
HAICOMM 0.80 0.89
@@
5.3 Analyses
We conduct a qualitative analysis of HAICOMM. In Figure 4, (a) and (b) are the input T1 and T2 MRIs, respectively. The table below shows the predictions by the three raters (Rater #1,#2,#3), then the predictions by SSR and ProMix trained with Rater #1’s labels and CROWDLAB’s labels (SSR w GT1, ProMix w GT1, SSR w CL GT, ProMix w CL GT). Next, we show SSR and ProMix trained with CROWDLAB’s labels and relying on human-AI collaborative classification (SSR w HAIC, ProMix w HAIC), followed by the result from our HAICOMM, and the ground truth label from surgical data. This figure shows that the proposed HAICOMM model can generate correct labels, with other approaches showing incorrect predictions. For (c) and (d), we show that the proposed HAICOMM model produced a correct label while most of other methods failed (only ProMix w GT1 and HAICOMM predict the surgical ground truth label correctly).
Figure 4: Qualitative analysis of HAICOMM. Each row (a),(b) and (c),(d) presents the input T1 and T2 MRI images, with corresponding tables below showing predictions from three human raters (Rater #1, #2, and #3). The tables also display predictions from SSR and ProMix models trained using labels from Rater #1 (SSR w/ GT1, ProMix w/ GT1) and CROWDLAB (SSR w/ CL GT, ProMix w/ CL GT). Following these, we provide the predictions from SSR and ProMix models trained on CROWDLAB labels and utilizing human-AI collaborative classification (SSR w/ HAIC, ProMix w/ HAIC). Finally, we present results from our HAICOMM model, along with the ground truth label based on surgical data.
In terms of the normal and abnormal volumes, they are sampled from distributions with equivalent condition / equipment / resolution, so they have similar properties. We calculate the statistics of the volume size distribution of the normal / abnormal (for the training data, we use majority vote results as the label) ones in Table 7 and Table 8. Moreover, we plot the distribution of manufacturer model of scanners for normal and abnormal in Figure 5 and the distribution of magnetic field strength for normal and abnormal in Figure 6.
Table 6: Number of normal and abnormal (i.e., with POD obliteration) volumes in the training and testing sets. In the training set, ‘A1’ to ‘A3’ are the three annotators, ‘MV’ represents the annotations from majority voting, and ‘CL’ denotes the annotations from CrowdLab [55]. In the testing set, we have the grount truth (GT) from surgery.
POD Training Set Testing Set
A1 A2 A3 MV CL GT
Normal 62 48 36 51 47 15
Abnormal 20 34 46 31 34 15
Table 7: Statistics of the volume size distribution of the T1 MRIs of POD obliteration dataset.
(a) Normal Statistics
Dim Mean Std Max Min
Depth 382.73 91.05 576 260
Height 313.67 74.57 464 260
Width 72.44 7.81 80 52
(b) Abnormal Statistics
Dim Mean Std Max Min
Depth 381.71 90.31 576 320
Height 309.35 71.95 464 250
Width 72.07 8.67 80 51
Table 8: Statistics of the volume size distribution of the T2 MRIs of POD obliteration dataset.
(a) Normal Statistics
Dim Mean Std Max Min
Depth 389.82 18.54 448 384
Height 190.06 17.64 256 176
Width 292.36 13.90 336 288
(b) Abnormal Statistics
Dim Mean Std Max Min
Depth 389.79 18.40 448 384
Height 193.05 28.50 336 160
Width 288.82 20.83 354 192
Figure 5: Distribution of manufacturer model of scanners for the normal (a) and abnormal (b) datasets.
Figure 6: Distribution of magnetic field strength of scanners for the normal (a) and abnormal (b) datasets.
@@ 6 Conclusion and Future Work
One potential limitation of our method is the dataset size. Currently, we are dedicated to collect more data from different clinical sources to expand the dataset. The use of such multiple clinical sources will require the exploration of domain adaptation techniques to enable a better flexibility of the method to work in multiple domains. Beyond this issue, the need for specific labellers for training and testing is another potential limitation. We plan to address this issue with the development of techniques that work with a variable set of labellers during training and testing. Another interesting direction is the collection of new datasets for other multi-modal multi-rater clinical problems to enable the evaluation of our HAICOMM in a different task. One more potential future direction is the development of a detector of region of interests, which will require dense (i.e., pixel-wise) annotations of the training and testing sets and the collection of significantly larger training sets.